OpenStudy (anonymous):

in DEF, DE = 15 and measurement of F = 32. find EF to the nearest tenth.

OpenStudy (dumbcow):

Are we assuming this is a right triangle?

OpenStudy (dumbcow):

If so, then sin 32 = 15/EF EF = 15/sin 32